### 22 Apr 2026 v2

# Kardashev’s Conundrum: Statistical Falsification of the Standard Kardashev Model and the Kardashev–Sagan–Nakamoto Resolution

Abstract
We test the standard Kardashev conjecture, which models the energy production of technologically advanced civilisations as growing at a fixed rate of one percent per year, against six decades of global energy production data (1965–2024) drawn from the Our World in Data energy dataset. The standard Kardashev one-percent exponential model is a poor fit to the data: the posterior growth rate obtained via Markov Chain Monte Carlo (MCMC) inference is $r=2.01\%\pm 0.03\%\,\mathrm{yr}^{-1}$ (95% credible interval $1.94\%,2.08\%]$), placing the Kardashev one-percent value well outside the credible interval. A linear Ordinary Least-Squares (OLS) model provides an excellent fit ($R^{2}=0.987$) and is preferred over the free-parameter exponential model by the Widely Applicable Information Criterion ($\Delta\mathrm{WAIC}=5.5$). The year-over-year energy increments exhibit significant non-Gaussian structure (Shapiro–Wilk $W=0.925$, $p=0.0014$), with negative skewness (skewness =0.664=-0.664) attributable to identifiable crisis events (2008, 2020), revealing path-dependence incompatible with the independent-increment multiplicative structure with positive drift required by Kardashev’s $(1+x)^{t}$ geometric series. However, extrapolation of the linear model to the Kardashev Type II threshold — the total bolometric luminosity of the Sun, $L_{\odot}=3.828\times 10^{26}\,\mathrm{W}$ — yields a civilisational transition timescale of $\sim 1.6\times 10^{15}$ yr: $\sim 10^{5}$ times both the age of the Universe and the main-sequence lifetime of the Sun. We demonstrate that no functional form fitted to energy production alone can simultaneously satisfy statistical adequacy and physical coherence, because the Kardashev state variable is dimensionally incomplete as a measure of civilisational advancement. To resolve the Conundrum, we propose the Kardashev–Sagan–Nakamoto (KSN) model, constructed by renormalising the energy production ordinate by the annual average Bitcoin network hashrate $\mathcal{H}(t)$, yielding the KSN state variable $\mathscr{B}(t)\;=\;P(t)/\mathcal{H}(t)$ in units of J Hash-1 — the KarNak unit. Any subsequent model fitted to $\mathscr{B}(t)$ carries the same parameter count as the corresponding model fitted to $P(t)$; the variable transformation itself consumes no degrees of freedom and is motivated by the Landauer limit connecting energy to irreversible computation, and fulfils the original requirement that the Kardashev scale be sensitive to the information richness of a civilisation. The KSN variable spans 14 orders of magnitude over 2009–2024, reflecting the dramatic improvement in the energy efficiency of Bitcoin proof-of-work computation; a formal functional fit to $\mathscr{B}(t)$ is deferred to a future work. We further discuss the implications of the KSN model for open astronomy development via decentralised blockchain protocols, and propose the KarNak unit as a universal information-energy yardstick for the Kardashev scale.
keywords:
Kardashev scale , Astroinformatics , Bayesian inference , Bitcoin , Blockchain , Open Development , Information theory , Civilization energy
1 Introduction
In a landmark 1964 paper, Kardashev (1964) proposed a universal scale for classifying the technological advancement of civilisations based on their total rate of energy consumption. On the Kardashev scale, a Type I civilisation harnesses all energy available on its home planet — at the time of writing, Kardashev (1964) estimated Earth’s total energy expenditure at $4\times 10^{19}$ erg s-1 ($4\times 10^{12}$ W) — a Type II civilisation consumes the total energy output of its host star ($L_{\odot}\approx 3.828\times 10^{26}$ W), and a Type III civilisation commands the energy output of its entire host galaxy. Kardashev (1964) noted that the annual increase in human energy expenditure was projected at 3–4% over the following 60 years (Putnam, 1948) on the basis of statistical findings. However, adopting a conservative growth rate of $x=1\%$ per year, Kardashev (1964) calculated that human civilisation would reach Type II status in approximately 3,200 years, and Type III status in approximately 5,800 years. This conjecture has since informed the search for extraterrestrial intelligence (Gray, 2020) and broader discussions of civilisational sustainability.
The notion that civilisational advancement is measured not only by energy throughput but by the efficiency with which energy is converted into organised information is ancient. The Antikythera mechanism, discovered in 1901 among the remains of a first-century BC shipwreck, is widely recognised as the earliest known analogue computational device (Freeth et al., 2021). Constructed from bronze gears with tolerances remarkable for their era, it calculated the positions of the Sun, Moon, and planets — converting mechanical energy into astronomical information with a sophistication not matched in Europe for over a millennium. From Babbage’s difference engine through Turing’s theoretical foundations of computation (Swade, 2017) to the Application-Specific Integrated Circuits (ASICs) that today secure the Bitcoin network, humanity has followed a continuous trajectory of reducing the energy cost per unit of computational work. The Kardashev–Sagan–Nakamoto model proposed here is the quantitative expression of this trajectory: it renormalises the Kardashev state variable by the energy cost per unit of proof-of-work computation, tracing civilisational advancement from the Antikythera mechanism to the present day.
Despite its influence, the Kardashev scale carries a fundamental limitation that has received little quantitative attention. Kardashev (1964), Sagan (1973), and Shklovskii and Sagan (1966) noted that an advanced civilisation must be characterised not only by its energy throughput but by the information richness of what it does with that energy. Sagan (1973) proposed that a civilisation’s state should also reflect its capacity to secure and transmit vast quantities of information, and suggested a binary-tree data structure as a measure of informational complexity. However, neither author formalised this requirement as a quantitative variable within the Kardashev framework. A civilisation that wastes energy inefficiently scores identically on the Kardashev scale to one that channels equivalent power into sophisticated computation. The Kardashev state variable — power in watts — is therefore dimensionally incomplete as a measure of civilisational advancement: it captures the quantity of energy consumed but not the quality of its use.
A second limitation emerges from direct comparison with data. When the standard one-percent exponential growth model of Kardashev (1964) is compared against six decades of observed total global energy production (Ritchie et al., 2022) — used here as a valid upper-limit proxy for the consumption metric of Kardashev (1964), following the argument that production necessarily equals or exceeds consumption at the civilisational scale and that the two track each other closely over decadal baselines — the model diverges substantially from the data. A linear Ordinary Least-Squares (OLS) model fits the same data exquisitely, but when extrapolated to the Kardashev Type II threshold it predicts a civilisational transition timescale of $\sim 1.2\times 10^{5}$ Hubble times: a physical reductio ad absurdum. We term this conflict Kardashev’s Conundrum: no functional form fitted to energy production alone can simultaneously satisfy statistical adequacy and physical coherence.
To resolve the Conundrum we propose the Kardashev–Sagan–Nakamoto (KSN) model, constructed by renormalising the energy production ordinate by the annual average Bitcoin network hashrate $H(t)$ (Nakamoto, 2009). The resulting state variable $\mathscr{B}=P/\mathcal{H}$, measured in units of J Hash-1 — which we term the KarNak unit — connects energy to irreversible computational work via the Landauer limit (Landauer, 1961), and fulfils the information-richness requirement identified but never quantified by Kardashev (1964) and Sagan (1973). The Bitcoin network is selected because it provides the only publicly auditable, unforgeable, and continuously updated global measure of proof-of-work computation (Nakamoto, 2009), requiring no free parameters beyond those of the original Kardashev model. As the KSN variable spans 14 orders of magnitude over 2009–2024, it traces the same trajectory of energy-to-computation efficiency that began with the Antikythera mechanism two millennia ago.
This paper is organised as follows. In Section 2.1 we describe the global energy production dataset and Bitcoin hashrate data. In Section 2.2 we assess the standard Kardashev model statistically. In Section 2.3 we present the linear OLS model and derive the Conundrum quantitatively. In Section 2.4 we perform Bayesian model comparison via Markov Chain Monte Carlo (MCMC) inference and the Widely Applicable Information Criterion (Watanabe, 2010). In Section 2.5 we introduce the KSN renormalization and its physical motivation. In Section 2.6 we summarise civilisational timescales on the KSN scale. In Section 2.7 we present the independence property argument as a further statistical objection to exponential growth models. In Section 3 we discuss falsifiability, limitations, and implications. Section 4 presents our conclusions.
The ancient Greek astronomers defined the logarithmic magnitude system and introduced concepts of limits and infinitesimals that would become the basis of differential geometry, calculus, and numerical methods. They also paved the way for the Julian Day — the counting yardstick used on the time axis of Figure 1 — which the Roman empire defined using the 28-year Solar, 19-year Metonic, and 15-year Indiction (tax) cycles. Many of these advances are found today in open-source toolkits for astronomy built by community efforts, e.g. SciPy, Astropy, and the IVOA standards.
It is now known after a plethora of multidisciplinary studies including that of Freeth et al. (2021) that during the ‘decline’ of Ancient Greece the first analogue ‘Computer’ was temporally ‘lost’ to most of the technological world, until when in the fourteenth century comparable ‘advanced’ mechanisms were developed and used in Astronomical clock machinery in Europe.
Roughly two-thousand years later, Charles Babbage and Alan Turing (AT) are widely recognised as the modern ’fathers’ of the digital computer (Swade, 2017). The pioneering work of AT paved the foundations of Artificial Intelligence (AI) and Cryptography.
2 The Kardashev Conundrum and the Kardashev–Sagan–Nakamoto Model
We now present our central analysis. We compare the standard Kardashev one-percent exponential model against a linear Ordinary Least-Squares (OLS) model fitted to the total global energy production data, and demonstrate that both models are unsatisfactory in complementary ways. We then propose the Kardashev–Sagan–Nakamoto (KSN) renormalization as the minimal resolution that is simultaneously consistent with the data and with physical causality.
2.1 Data
The energy production data $\mathcal{D}$ consist of the World total primary energy consumption in terawatt-hours per year (TWh yr-1), drawn from the Our World in Data (OWID) energy dataset (Ritchie et al., 2022), which integrates the Energy Institute (EI) Statistical Review of World Energy and the US Energy Information Administration (EIA) international energy series; see Smil (2017) for a comprehensive historical analysis of the industrial-era energy trajectory. We use $N=60$ annual data points spanning 1965–2024. The OWID dataset column primary_energy_consumption reports values using the direct primary energy accounting method, which measures the physical energy output of each source without applying fossil-fuel-equivalent conversion factors. This is distinct from the substitution method historically used by the Energy Institute, which inflates non-fossil contributions (nuclear, hydro, wind, solar) by a thermal efficiency factor ($\sim 0.38$$0.40$) to estimate the fossil fuel input that would have produced the same electricity. We adopt the direct method throughout because it measures the actual rate of energy flow through civilisation — the physical quantity that the Kardashev scale is defined in terms of — and avoids an accounting convention whose magnitude grows with the renewable energy share. We verify in Section 3.2 that all qualitative conclusions hold under the substitution method.
All energy values are converted to SI watts (J s-1) assuming uniform power delivery over each calendar year via:
$P\,[\mathrm{W}]\;=\;E\,[\mathrm{TWh\,yr}^{-1}]\;\times\;\frac{10^{12}\,[\mathrm{Wh\,TWh}^{-1}]\times 3600\,[\mathrm{J\,Wh}^{-1}]}{365.25\times 24\times 3600\,[\mathrm{s\,yr}^{-1}]}\,$, (1)
where the numerator converts terawatt-hours to joules ($1\,\mathrm{TWh}=3.6\times 10^{15}\,\mathrm{J}$) and the denominator is the number of seconds in one Julian year. The factors of $3600$ cancel, reducing Equation (1) to $P=E\times 10^{12}/(365.25\times 24)$ W. This yields $P(1965)=4.95\,\mathrm{TW}$, in close agreement with the Kardashev (1964) baseline of $4\,\mathrm{TW}$, and $P(2024)=20.16\,\mathrm{TW}$, consistent with the International Energy Agency’s direct-method estimate of $\sim 20\,\mathrm{TW}$.
We set the origin of the time coordinate at $t_{0}=1964$, the year of the Kardashev conjecture (Kardashev, 1964), so that $t=\mathrm{year}-1964$ throughout.
The Bitcoin network hashrate data $\mathcal{H}(t)$ are yearly averages in hash s-1 for 2009–2024. These span the full history of Bitcoin mining from the CPU era ($\sim 7\,\mathrm{MH\,s^{-1}}$ in 2009) through the ASIC era ($\sim 6.2\times 10^{20}\,\mathrm{H\,s^{-1}}$ in 2024), covering 16 annual data points and 14 orders of magnitude in hashrate. For 2009–2013 (the CPU and GPU eras), no single authoritative source of annual-average hashrate exists; we derive yearly averages from the on-chain difficulty record using the standard relation $\mathcal{H}=D\times 2^{32}/600$, taking the logarithmic mean of the January 1 and December 31 difficulty values to obtain a time-weighted annual average appropriate for an exponentially growing quantity. For 2014–2024 (the ASIC era), we use the Blockchain.com and NASDAQ BCHAIN/HRATE annual averages directly.
2.2 The Standard Kardashev Model and its Statistical Assessment
Kardashev (1964) proposed that the energy production of a technologically advanced civilisation increases at a constant fractional rate per year, giving a geometric series:
$P_{\rm K}(t)\;=\;P_{0}\,(1+x)^{t}\;\approx\;P_{0}\,e^{xt}\,,\quad x=0.01\,$, (2)
where $P_{0}$ is the energy production at $t=0$ (1964) and $x=0.01$ corresponds to the stated one-percent per year growth rate. Using this model, Kardashev (1964) calculated that a Type II civilisation — defined as one harnessing the total bolometric luminosity of its host star, $L_{\odot}=3.828\times 10^{26}\,\mathrm{W}$ (Mamajek et al., 2015) — would be reached in approximately 3,2003{,}200 years from 1964, corresponding to approximately year 5188 CE. Sagan (1973) obtained a similar value using a slightly revised Type II threshold of $\sim 10^{26}\,\mathrm{W}$.
Figure 1 shows the Kardashev model (Eq. 2) alongside the OWID data, together with the Type I and Type II civilisation limits and the solar insolation at Earth ($F_{\oplus}\approx 1.74\times 10^{17}\,\mathrm{W}$). The one-percent Kardashev model diverges visibly from the data over the full six-decade baseline.
We quantify this divergence statistically. The likelihood of the data given the Kardashev model with a point prior on $x=0.01$ is:
$\mathcal{L}(\mathcal{D}\mid\theta_{\rm K},M_{\rm K})\;=\;\prod_{i=1}^{N}\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\!\left[-\frac{\bigl(P_{i}-P_{0}\,(1.01)^{t_{i}}\bigr)^{2}}{2\sigma^{2}}\right],$ (3)
where $\sigma^{2}$ is estimated from the residuals. The residual sum of squares for the Kardashev model exceeds that of the linear model by more than two orders of magnitude, constituting overwhelming evidence against the one-percent prior. We therefore conclude that the standard Kardashev exponential model with $x=0.01$ is statistically untenable.
2.3 The Linear Model: an Excellent Fit with a Physical Reductio
We next consider a linear OLS model:
$P_{\rm L}(t)\;=\;a+b\,t\,$, (4)
with parameters $a$ and $b$ estimated by minimising the sum of squared residuals. Following Occam’s razor (Blumer et al., 1987), this two-parameter model is the simplest physically plausible alternative. The OLS solution gives $b=2.44\times 10^{11}\,\mathrm{W\,yr}^{-1}$ and coefficient of determination $R^{2}=0.987$. The fit is excellent over the six-decade baseline of the data (Figure 1). Figure 2 shows the same data and fits on a linear vertical axis, making the remarkably low dispersion directly visible: the linear and exponential models are visually near-identical across 1965–2024, while the Kardashev one-percent model diverges strikingly downward from the data.
To assess the normality of the residual structure, we compute the year-over-year absolute differences $\Delta P_{i}=P_{i}-P_{i-1}$ and apply the Shapiro–Wilk normality test (Shapiro and Wilk, 1965). We obtain $W=0.925$, $p=0.0014$: we reject the null hypothesis that $\Delta P$ is drawn from a Normal distribution. The distribution of $\Delta P$ (Figure 3) exhibits significant negative skewness (skewness =0.664=-0.664), driven by identifiable exogenous shocks to global energy production at 2008 (subprime financial crisis) and 2020 (COVID-19 pandemic). This non-Gaussian structure constitutes direct evidence of path-dependence in the global energy system: large negative shocks are systematically more probable than an independent-increment multiplicative generator with positive drift would predict, and their occurrence is traceable to specific historical events rather than symmetric statistical fluctuations.
A further structural objection to $(1+x)^{t}$ with $x>0$ is that the model is sign-constrained: every year must exhibit positive growth by construction. The observed series contains identifiable years of negative growth (e.g., 2008 and 2020) that the model cannot accommodate at any parameter value — a structural, not merely statistical, falsification of the always-positive growth assumption that is independent of the non-Gaussianity and path-dependence arguments above.
Notwithstanding its excellent in-sample fit, the linear model fails catastrophically on extrapolation. Solving $P_{\rm L}(t^{*})=L_{\odot}$ gives:
$t^{*}\;=\;\frac{L_{\odot}-a}{b}\;\approx\;1.6\times 10^{15}\,\mathrm{yr}\;\approx\;1.2\times 10^{5}\,H_{0}^{-1}\,$, (5)
where $H_{0}^{-1}\approx 13.8\,\mathrm{Gyr}$ is the Hubble time. The implied Type II timescale exceeds the age of the Universe by five orders of magnitude. This is not merely improbable — it is physically incoherent, since it vastly exceeds the main-sequence lifetime of the Sun and the projected heat death of the Universe. We designate this result Kardashev’s Conundrum: the linear model, which best fits the data by all standard statistical criteria, predicts a civilisational transition timescale that is cosmologically nonsensical.
2.4 Model Comparison via MCMC and WAIC
To place the comparison on a rigorous Bayesian footing, we fit an exponential model with the growth rate $r$ as a free parameter:
$P_{\rm E}(t)\;=\;a_{0}\,e^{rt}\,$, (6)
using a Metropolis–Hastings Markov Chain Monte Carlo (MCMC) sampler (Metropolis et al., 1953; Hastings, 1970) with a prior $r\sim\mathcal{N}(0.01,\,0.02^{2})$ centred on the Kardashev value and N=60, [^0] N=60{,}000 posterior samples after a 15,00015{,}000-step burn-in. The burn-in phase is discarded because the first iterations of the chain are influenced by the arbitrary starting point in parameter space rather than by the target posterior; the Markov property of the sampler ensures that the chain eventually forgets this initial condition and converges to the stationary distribution, from which the retained samples are drawn. The posterior mean growth rate is $r=2.01\%\pm 0.03\%\,\mathrm{yr}^{-1}$ (95% credible interval $1.94\%,2.08\%]$), with the Kardashev value of $1\%\,\mathrm{yr}^{-1}$ lying well outside the $95\%$ credible interval. The best-fit exponential model gives $R^{2}=0.987$, nearly identical to the linear model ($\Delta R^{2}<0.001$).
The Widely Applicable Information Criterion (WAIC; Watanabe, 2010) comparison yields $\Delta\mathrm{WAIC}=\mathrm{WAIC}_{\rm exp}-\mathrm{WAIC}_{\rm linear}=5.5$: the linear model is preferred over the free-$r$ exponential. Since both models have $K=2$ free parameters, the AIC and BIC penalties are identical, and the WAIC preference for the linear model reflects genuine differences in predictive accuracy. Together, the near-identical $R^{2}$ values and positive $\Delta\mathrm{WAIC}$ confirm the statistical statement of Kardashev’s Conundrum: the data cannot distinguish between the two functional forms on the observed baseline, and of the two, the simpler linear model is marginally preferred.
The complete statistical picture is:
  1. 1.
    The one-percent Kardashev exponential is falsified by the data (poor fit; posterior $r=2.01\%$, well outside the one-percent prior).
  2. 2.
    The linear model and the free-$r$ exponential have statistically identical in-sample performance ($\Delta R^{2}<0.001$), with the linear model marginally preferred ($\Delta\mathrm{WAIC}=5.5$).
  3. 3.
    The year-over-year differences $\Delta P$ are non-Gaussian (Shapiro–Wilk $W=0.925$, $p=0.0014$), exhibiting path-dependent negative skewness inconsistent with any independent-increment multiplicative generator with positive drift.
  4. 4.
    The linear model is physically falsified by extrapolation (Type II timescale $\sim 1.2\times 10^{5}\,H_{0}^{-1}$).
No functional form fitted to $P(t)$ can simultaneously satisfy statistical adequacy and physical coherence. This is the Conundrum, stated precisely.
2.5 The KSN Resolution: Renormalization to J Hash-1
The resolution of Kardashev’s Conundrum cannot be achieved by selecting a different fitting function for the same state variable $P(t)$, because the problem is not one of functional form — it is one of variable choice. We demonstrate this via the following argument.
Dimensional incompleteness. The Kardashev scale measures energy throughput, $P$ [W = J s-1], but makes no reference to what that energy accomplishes. As Sagan (1973) noted, a civilisation that wastes energy inefficiently is rated equally to one that channels the same power into information-rich computation. The scale is therefore dimensionally incomplete as a measure of civilisational advancement: it conflates energy expenditure with technological capability.
The Landauer connection. The fundamental physical link between energy and information is the Landauer limit (Landauer, 1961; Bennett, 1982): the minimum energy required to erase one bit of information at temperature $T$ is $k_{B}T\ln 2\approx 2.85\times 10^{-21}\,\mathrm{J}$ at room temperature. The ratio $P/H$ — global energy production divided by global computational hashrate — has dimensions of energy per hash, and is an empirical measure of how efficiently a civilisation converts energy into irreversible computational work. As $P/H\to k_{B}T\ln 2$ (per bit), a civilisation approaches the thermodynamic limit of computation. Our physical variable will thus inherit physical meaning absent from $P(t)$ alone.
The renormalization. We define the KSN state variable as:
$\mathscr{B}(t)\;=\;\frac{P(t)}{\mathcal{H}(t)}\qquad\left[\mathrm{J\,Hash^{-1}}\equiv\mathrm{KN}\right]\,$, (7)
where $\mathcal{H}(t)$ is the yearly average Bitcoin network hashrate in Hash s-1. We use the Bitcoin hashrate because it is the only publicly auditable, unforgeable, and continuously updated global measure of computational proof-of-work (Nakamoto, 2009), fulfilling the requirement of Sagan (1973) that the Kardashev state variable be sensitive to the “security of vast quantities of information.” No free parameter is added to the model: the renormalization is a ratio of two independently measured physical quantities.
Statistical motivation. The variable transformation $P\to\mathscr{B}=P/\mathcal{H}$ is not a Box–Cox power transformation of $P$ (Box and Cox, 1964), which would remain within the family of functional forms fitted to the original variable. It is instead a ratio normalization by an external physically motivated quantity, taking the analysis outside the Box–Cox family. The KSN variable $\mathscr{B}(t)$. is our proposal for the correct extensive variable of the Kardashev scale.
Concurrent work. Complementary work from the quantum-computing perspective has recently examined Bitcoin mining as a Kardashev-scale problem (Dallaire-Demers and BTQ Technologies Team, 2026), demonstrating that quantum-accelerated Bitcoin mining at current network difficulty would require a fleet of $\sim 10^{23}$ physical qubits drawing $\sim 10^{25}\,\mathrm{W}$ — resources consistent with a Kardashev Type II civilisation. Their result places an upper bound on quantum-mining feasibility: the resource requirements are so extreme that practical quantum mining collapses under hardware, error-correction, and fleet-logistics overheads. Our analysis is methodologically distinct — we use the Bitcoin hashrate as an empirical proxy for global classical proof-of-work and renormalise the Kardashev state variable accordingly — but shares the use of proof-of-work infrastructure as a natural bridge between civilisational energy and irreversible computation.
Figure 4 shows $\mathscr{B}(t)$ for 2009–2024, spanning 16 annual data points and 14 orders of magnitude in $\mathscr{B}$. The data display a monotonically decreasing trend reflecting the rapid improvement in Bitcoin proof-of-work energy efficiency from the CPU era (2009) through the GPU era (2010–2012) to the ASIC era (2013–2024). The structural breaks between these eras preclude a single-exponential description of $\mathscr{B}(t)$; a formal era-segmented functional fit is deferred to a future paper.
2.6 Civilisational Timescales on the KSN Scale
Table 1 summarises the Type II timescales implied by each model. The Kardashev and Sagan exponential models give timescales of order $10^{3}\,\mathrm{yr}$. These are reached tautologically: any exponential model of the form $P=P_{0}e^{rt}$ will reach any finite threshold at some finite time, so the Type II crossing year is determined by the assumed growth rate $r$ and is not a prediction in any falsifiable sense. The linear model, as shown in Eq. (5), gives a timescale that is cosmologically absurd. On the KSN scale, the concept of a Type II threshold must be re-expressed in terms of $\mathscr{B}=P/\mathcal{H}$: a Type II KSN civilisation is one for which $\mathscr{B}$ approaches the Landauer limit, i.e. the energy cost per hash approaches $k_{B}T\ln 2$. This reframing is consistent with the original intent of Kardashev (1964) and the information-theoretic extension proposed by Sagan (1973).
Table 1: Type II civilisation timescales under each model. All timescales are measured from 1964. $H_{0}^{-1}=13.8\,\mathrm{Gyr}$ is the Hubble time. The MCMC exponential uses the posterior mean $r=2.01\%\,\mathrm{yr}^{-1}$.
Model Growth rate Type II year (CE) $\Delta t/H_{0}^{-1}$ Kardashev 1964 (fixed $r$) $1.00\%\,\mathrm{yr}^{-1}$ $\sim 5188$ $\sim 2\times 10^{-7}$ Exp. free $r$ (MCMC mean) $2.01\%\,\mathrm{yr}^{-1}$ $\sim 3547$ $\sim 1\times 10^{-7}$ Linear OLS $2.44\times 10^{11}\,\mathrm{W\,yr}^{-1}$ $\sim 1.6\times 10^{15}$ $\sim 1.2\times 10^{5}$ KSN ($\mathscr{B}=P/\mathcal{H}$) $\mathscr{B}\to k_{B}T\ln 2$ Re-expressed via Landauer limit
2.7 The Independence Property as a Further Objection to the Exponential
A complementary statistical objection to the exponential Kardashev model arises from probability theory. Kardashev’s geometric series $(1+x)^{t}$ implicitly models year-over-year growth factors as independent draws from a distribution with mean $(1+x)$. For a non-negative random variable $X$ with this property, the memoryless relation holds:
$P(X>a+b\mid X>b)\;=\;P(X>a)\,.$ (8)
That is: the fractional increment $\Delta P/P$ is independent of the current level and of history. The global energy system, however, manifestly possesses memory: infrastructure lock-in, geopolitical constraints, technology diffusion rates, and crisis events (e.g., 2008, 2020) all induce path-dependence in the year-over-year increments, violating the independence assumption that $(1+x)^{t}$ requires.
The Shapiro–Wilk result ($W=0.925$, $p=0.0014$) provides direct empirical evidence of this path-dependence. The distribution of $\Delta P$ is non-Gaussian and negatively skewed, with the two largest negative outliers corresponding precisely to the 2008 global financial crisis and the 2020 COVID-19 pandemic. These are identifiable, historically specific events — not symmetric statistical fluctuations drawn from an independence-preserving distribution. An independent-increment multiplicative generator with positive drift, which $(1+x)^{t}$ instantiates, cannot accommodate large negative shocks driven by specific historical contingencies. The data refute this assumption. The independence property is therefore not just a mathematical abstraction: it is a testable claim about the generative process, and the evidence is against it.
Table 2: Comparison of energy production baselines and Type II civilisation timescale predictions. $P_{0}$ is the average continuous power at the start of each analysis, converted from the reported annual energy via Eq. (1). $H_{0}^{-1}=13.8\,\mathrm{Gyr}$ is the Hubble time.
Source $P_{0}$ Type II timescale Kardashev (1964) (1964 estimate) $4.00\,\mathrm{TW}$ $\sim 3.2\,\mathrm{kyr}$ at $r=1\%\,\mathrm{yr}^{-1}$ OWID (Ritchie et al., 2022) (1965, this work) $4.95\,\mathrm{TW}$ $\sim 1.2\times 10^{5}\,H_{0}^{-1}$ (linear OLS)
Figure 1: Global energy production 1965–2024 (OWID dataset) in watts, with the standard Kardashev one-percent exponential model (dashed-orange), the best-fit linear OLS model (solid-red), and the Kardashev Type I and Type II civilisation thresholds (Horizontal lines). The Kardashev model diverge visibly from the OLS data fit over the full six-decade time baseline.
Figure 2: Zoomed view of the OWID global primary-energy production data (1965–2024) with the linear and exponential best-fit models overlaid on a linear vertical axis (W), making visible the remarkably low dispersion of the data about both fitted models that is compressed in the logarithmic-axis view of Figure 1. The linear OLS fit (blue dashed, $R^{2}=0.987$) and the free-rate exponential fit ($r=2.01\%\,\mathrm{yr}^{-1}$; solid red, $R^{2}=0.987$) are visually near-identical over the six-decade baseline, confirming the WAIC result that the two functional forms are statistically indistinguishable on the observed data. The Kardashev one-percent exponential (salmon dotted) diverges dramatically downward from the data, falsifying the conservative assumption universally adopted in the subsequent literature. The Kardashev Type I threshold from 1964 (yellow dotted, $4\times 10^{12}\,\mathrm{W}$) lies well below the full dataset: Earth has long since surpassed Kardashev’s original Type I energy benchmark.
Figure 3: Year-over-year differences $\Delta P_{i}=P_{i}-P_{i-1}$ in global primary-energy production (W) as a function of calendar year 1965–2024, derived from the OWID direct-method energy dataset. Two stray negative outliers — clearly separated from the general trend of the data — occur at 2008 and $20$, corresponding respectively to the subprime great financial crisis and the COVID-19 pandemic. These are identifiable historical shocks, not symmetric statistical fluctuations. The Shapiro–Wilk test on the $\Delta P$ series yields $W=0.925$, $p=0.0014$, with skewness =0.664=-0.664. The distribution is significantly non-Gaussian, and the identifiable crisis outliers constitute direct evidence of path-dependence — incompatible with the structural assumptions of the Kardashev series $(1+x)^{t}$ (independent increments with positive drift).
Figure 4: The KSN state variable $\mathscr{B}(t)=P(t)/\mathcal{H}(t)$ in KarNak units (J Hash-1) for 2009–2024, spanning 14 orders of magnitude. The monotonically decreasing trend reflects the rapid improvement in Bitcoin proof-of-work energy efficiency from the CPU era (2009) through the GPU era (2010–2012) to the ASIC era (2013–2024). Structural breaks between eras preclude a single-exponential description; a formal fit is deferred to a future paper.
3 Discussion
3.1 Falsifiability of the KSN Model
A scientific model is meaningful only insofar as it generates predictions that can in principle be refuted by data (Blumer et al., 1987). We identify four concrete falsification criteria for the KSN model.
First, the posterior growth rate $r=2.01\%\pm 0.03\%\,\mathrm{yr}^{-1}$ is a quantitative prediction: if sustained observation of global energy production over the coming decades yields a growth rate consistently outside the 95% credible interval $1.94\%,2.08\%]$, the exponential component of the KSN framework is falsified.
Second, the non-Gaussian structure of $\Delta P$ (Shapiro–Wilk $W=0.925$, $p=0.0014$) establishes that year-over-year energy increments exhibit path-dependent behaviour driven by identifiable historical events. This is a testable claim: if future decades produce a $\Delta P$ distribution consistent with an independent-increment multiplicative generator with positive drift, the path-dependence argument is falsified.
Third, the KSN variable $\mathscr{B}=P/\mathcal{H}$ predicts a monotonically decreasing trend as Bitcoin ASIC efficiency improves toward the Landauer limit (Landauer, 1961). If the Bitcoin hashrate growth stalls or reverses permanently — for example if proof-of-work mining becomes economically unviable — the KSN variable would cease to decrease and the renormalization would lose its physical motivation. This is a genuine falsification risk that the model acknowledges explicitly.
Fourth, the dimensional incompleteness argument predicts that no functional form fitted to $P(t)$ alone can resolve Kardashev’s Conundrum. If a future author produces a model of $P(t)$ without additional state variables that gives both a statistically adequate fit to the data and a physically coherent Type II timescale, the KSN renormalization is unnecessary and the argument is falsified.
3.2 Limitations
Several limitations of the present analysis deserve explicit acknowledgement.
A methodological choice concerns the energy accounting convention. The OWID dataset column primary_energy_consumption reports values using the direct primary energy method, which records the physical energy output of each source. The alternative substitution method inflates non-fossil contributions (nuclear, hydro, wind, solar) by a thermal efficiency factor ($\sim 0.38$$0.40$) to estimate the fossil fuel input that would have produced the same electricity. We adopt the direct method because the Kardashev scale measures physical power, not fossil-fuel equivalents. We have verified that all qualitative conclusions — falsification of the one-percent model, preference for the linear fit, and the persistence of the Conundrum — hold under the substitution method, with quantitative differences: the substitution method yields $P(1965)=7.66\,\mathrm{TW}$, $b=3.12\times 10^{11}\,\mathrm{W\,yr}^{-1}$, $r=1.85\%\,\mathrm{yr}^{-1}$, and a Type II timescale of $\sim 8.9\times 10^{4}\,H_{0}^{-1}$. The choice of accounting method does not affect the dimensional incompleteness argument that motivates the KSN renormalisation. We note that the distinction between these methods will grow in significance as the global energy mix shifts toward nuclear and renewable sources — a transition that appears to be accelerating, with several nations expanding nuclear capacity — since the substitution method increasingly overstates the total energy budget relative to the physical power actually delivered.
The Bitcoin network hashrate is the only publicly auditable global proof-of-work dataset currently available, but it represents a single network. A more complete KSN state variable would integrate all global computational proof-of-work, including other proof-of-work cryptocurrencies and, eventually, any future large-scale computational infrastructure that provides an unforgeable public record. The present analysis should therefore be understood as a proof of concept using the best available proxy, rather than a definitive measurement of civilisational computational efficiency.
The hashrate data span only 2009–2024 — sixteen annual data points covering approximately one human generation. The KSN variable exhibits structural breaks between the CPU/GPU era (2009–2012) and the ASIC era (2013–2024), reflecting qualitative changes in the technology of proof-of-work computation rather than smooth continuous evolution. A formal era-segmented fit to $\mathcal{H}(t)$ is deferred to a future paper.
A further limitation concerns the production–consumption distinction. The OWID dataset reports total primary energy production, whereas Kardashev (1964) defined his scale in terms of energy consumption. We have argued that production constitutes a valid upper-limit proxy for consumption at the civilisational scale, and that the two track each other closely over decadal baselines. However, the ratio of production to consumption is not precisely unity and varies with time, introducing a systematic uncertainty that is difficult to quantify without a separate long-baseline consumption dataset. We note that our pull request to the OWID energy repository (Ritchie et al., 2022) identified and documented this label inconsistency between the column name primary_energy_consumption and the actual measured quantity, which is production.
3.3 The Novel Observation on Kardashev’s Growth Rates
[…]
3.4 Connection to the Drake Equation
The Drake equation for the number of communicating civilisations in the Galaxy includes a longevity factor $L$ — the average duration for which a civilisation remains detectable. Monte Carlo approaches to estimating the probability of causal contact between communicating civilisations (Lares et al., 2020) provide a complementary framework in which the KSN variable may eventually constrain the longevity factor. This factor carries the largest relative uncertainty of any term in the Drake equation and is sensitive to whether civilisations tend to self-destruct or achieve long-term stability (Gray, 2020).
The KSN model is relevant to the longevity factor in the following sense: a civilisation that approaches the Landauer limit — that is, one for which $\mathscr{B}\to k_{B}T\ln 2$ per bit — has achieved near-optimal use of available energy for computation, which we argue is a necessary condition for long-term civilisational stability. The KSN Type II threshold, defined as $\mathscr{B}$ approaching the Landauer limit, is therefore not merely an energy milestone but an information-thermodynamic one. Whether $L$ is large or small may depend on whether civilisations reach this threshold before exhausting or destabilising their energy resources — a question the KSN model frames quantitatively for the first time.
3.5 The Doomsday Clock
The Bulletin of the Atomic Scientists advanced the Doomsday Clock to 85 seconds to midnight in 2026 — the closest it has been to civilisational catastrophe in the history of the index. The KSN model provides an independent quantitative perspective on civilisational risk: the dramatic decrease in $\mathcal{H}(t)$ observed over 2009–2024 reflects an exponential improvement in the energy efficiency of global computation, consistent with a civilisation moving toward the Landauer limit. Whether this trajectory can be sustained, and whether it will be accompanied by the institutional stability required to avoid catastrophe, lies beyond the scope of this model — but the KSN framework provides the quantitative language in which such questions can eventually be posed. We note that in his 1985 congressional testimony on climate change, Sagan (1985) identified the need for “a degree of international amity which certainly doesn’t exist today.” We speculate that decentralised proof-of-work protocols, by providing trustless cooperation without requiring political goodwill, may offer a partial technological foundation for the amity Sagan envisioned.
4 Conclusions
We have tested the standard Kardashev one-percent exponential model against six decades of global energy production data and found it statistically untenable. The posterior growth rate from MCMC inference is $r=2.01\%\pm 0.03\%\,\mathrm{yr}^{-1}$ (95% credible interval $1.94\%,2.08\%]$), placing the Kardashev one-percent value well outside the credible interval. The year-over-year energy differences exhibit significant non-Gaussian, path-dependent structure (Shapiro–Wilk $W=0.925$, $p=0.0014$), with negative skewness attributable to identifiable crisis events (2008, 2020), incompatible with the independent-increment multiplicative structure with positive drift required by Kardashev’s $(1+x)^{t}$ geometric series.
A linear OLS model fits the data excellently ($R^{2}=0.987$) and is marginally preferred over the free-parameter exponential model by the Widely Applicable Information Criterion ($\Delta\mathrm{WAIC}=5.5$). However, extrapolation of the linear model to the Kardashev Type II threshold yields a civilisational transition timescale of $\sim 1.2\times 10^{5}$ Hubble times — a physical reductio ad absurdum we term Kardashev’s Conundrum. Long before this date, the Sun will have exhausted its hydrogen fuel, expanded into a red giant, and likely engulfed or scorched the Earth. No functional form fitted to energy production alone can simultaneously satisfy statistical adequacy and physical coherence, because the Kardashev state variable is dimensionally incomplete as a measure of civilisational advancement.
We have proposed the Kardashev–Sagan–Nakamoto (KSN) model as a resolution: renormalising the energy production ordinate by the annual average Bitcoin network hashrate $\mathcal{H}(t)$, yielding the KSN state variable $\mathscr{B}(t)\;=\;\frac{P(t)}{\mathcal{H}(t)}$ in units of J Hash-1 — the KarNak unit. This renormalisation adds no free parameters, is physically motivated by the Landauer limit connecting energy to irreversible computation (Landauer, 1961), and fulfils the requirement of Sagan (1973) that the Kardashev scale be sensitive to the information richness of a civilisation. The KSN variable spans 14 orders of magnitude over 2009–2024, tracing the trajectory of energy-to-computation efficiency from the CPU era through to modern ASICs.
[…]
A formal era-segmented functional fit to $\mathcal{H}(t)$, accounting for the structural breaks between the CPU, GPU, and ASIC eras of Bitcoin mining, is deferred to a future work. The KarNak unit is proposed as a universal information-energy yardstick for the Kardashev scale, and the KSN framework as a quantitative language for discussing civilisational longevity in the context of the Drake equation.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author used Claude (Anthropic) in order to assist with statistical verification, code debugging, dimensional analysis, and manuscript editing. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
5 Acknowledgments
SG thanks R. Gamen from the VO of Argentina (NOVA) for the opportunity to serve as Argentina’s representative on the IVOA Executive Committee (2015–2024). SG also thanks colleagues at the IATE and the OAC for their support. This work made use of Astropy, a community-developed core Python package for Astronomy (Robitaille et al., 2013), and the broader open-source scientific software ecosystem stack including SciPy and NumPy.Simulations were run on the Sersic cluster at the ‘Instituto de Astronomía Teórica y Experimental’ (IATE).This research was partially conducted during a period of unpaid leave from CONICET, the Astronomical Observatory of Córdoba and the National University of Córdoba, motivated by national government austerity measures that suspended research funding across Argentine scientific institutions.
Author contributions (CRediT): S. Gurovich: Conceptualization, Methodology, Software, Formal analysis, Investigation, Writing – original draft, Writing – review & editing, Visualization, Project administration.